OpenStudy (anonymous):

factorize: a2+b2-2(ab-ac+bc)

mathslover (mathslover):

check it again @basith is it : \[\large{a^2+b^2+c^2-2(ab-ac+bc)}\] \[\large{a^2+b^2-c^2-2(ab-ac+bc)}\]

mathslover (mathslover):

which one?

mathslover (mathslover):

\[{\large(a+b+c)^2 = a^2+b^2+c^2+2(ab+bc+ca)}\]

OpenStudy (anonymous):

\[a ^{2}+b ^{2}-2(ab-ac+bc)\]

mathslover (mathslover):

ok .. : \[{a^2+b^2-2(ab-ac+bc)=a^2+b^2-2ab+2ac-2bc}\] \[\large{=a^2-2ab+b^2-2bc+2ac}\]

mathslover (mathslover):

\[\large{a(a-2b)+b(b-2c)+2ac}\]

OpenStudy (neo92):

a^2 - 2ab + b^2 -2bc + 2ac = (a-b)^2 + c(a-b) = (a-b)(a-b+c)

OpenStudy (neo92):

got it??

mathslover (mathslover):

\[\large{a^2-2ab+b^2-2bc+2ac}\] \[\large{(a-b)^2+2c(a-b)}\] \[\large{(a-b)(a-b+2c)}\] good work @Neo92

mathslover (mathslover):

@basith got it?

OpenStudy (neo92):

@ mathsolver thnx.. i missed da 2 there..

mathslover (mathslover):

no problem ..

OpenStudy (anonymous):

i have a doubt @mathslover

mathslover (mathslover):

ask @basith

OpenStudy (anonymous):

how did u get..... \[(a-b)(a-b+2c)\]

mathslover (mathslover):

i took (a-b) common

mathslover (mathslover):

i was having: \[\large{(a-b)^2+2c(a-b)}\] \[\large{(a-b)(a-b)+2c(a-b)}\] \[\large{(a-b)(a-b+2c)}\]

OpenStudy (anonymous):

(a-b)(a-b+2c) u can see http://www.wolframalpha.com/input/?i=factorize%3A+a2%2Bb2-2%28ab-ac%2Bbc%29